OpenStudy (anonymous):

i need to put 202 over 256 in radical expression in slimplest form

OpenStudy (e.mccormick):

OK. So start factoring them. Every time two factors match, you can putt them out. The obvious start is with powers of 2 because they are even. As a bigger hint, 256 is \(2^{something}\) so it will simplify a lot!

OpenStudy (e.mccormick):

putt... doh! Pull. As in \(\sqrt{20}=\sqrt{5\times 2\times 2}=2\sqrt{5}\) where I pulled out a pair of 2s.

OpenStudy (anonymous):

wait dpnt the factors have to be a prefect square

OpenStudy (anonymous):

i no its something over 16

OpenStudy (e.mccormick):

For a factor to come out, it needs to be a square. But with the root 20 example, you can see how some came out and some stayed in.

OpenStudy (anonymous):

so would 2 times 5 squared go over 16 (my computer wont let me put the radical sign)

OpenStudy (anonymous):

but wait its 202 not 20 were did u get 20

OpenStudy (e.mccormick):

I jsut used 20 as an example. 202 is 101*2, and 101 is a prime, so it just does not go anywhere and stays as the root of 202.

OpenStudy (e.mccormick):

So you actually knew the answer:\[\frac{\sqrt{202}}{16}\]

OpenStudy (anonymous):

oh ok thx its some much clearere now

OpenStudy (e.mccormick):

Yah, I should have said "As an example" rahter than just "as in" Might have confused you less.